OpenStudy (anonymous):

What is the equation of the line tangent to the curve 2x^2-y^4 = 1 at the point (-1,1)?

OpenStudy (experimentx):

use implicit differentiation to find slope at that point. then calculate the tangent using that point.

OpenStudy (amistre64):

y=mx-mPx+Py ; given m=f'(Px) and the point (Px,Py)

OpenStudy (experimentx):

2x^2-y^4 = 1 \[ \frac{d}{dx}(2x^2-y^4) = d(1)/dx\] \[ => 4x - 4y^3 dy/dx = 0 \] \[ => dy/dx = x/y^3 = -1\] or slope = -1, If you know single point formula or slope intercept form then you will be able to find slope.